# ROVR-Open-Dataset: A Large-Scale Depth Dataset for Autonomous Driving

### v1

@@ Abstract
Depth estimation is a fundamental task for 3D scene understanding in autonomous driving, robotics, and augmented reality. Existing depth datasets, such as KITTI, nuScenes, and DDAD, have advanced the field but suffer from limitations in diversity and scalability. As benchmark performance on these datasets approaches saturation, there is an increasing need for a new generation of large-scale, diverse, and cost-efficient datasets to support the era of foundation models and multi-modal learning. To address these challenges, we introduce a large-scale, diverse, frame-wise continuous dataset for depth estimation in dynamic outdoor driving environments, comprising 20K video frames to evaluate existing methods. Our lightweight acquisition pipeline ensures broad scene coverage at low cost, while sparse yet statistically sufficient ground truth enables robust training. Compared to existing datasets, ours presents greater diversity in driving scenarios and lower depth density, creating new challenges for generalization. Benchmark experiments with standard monocular depth estimation models validate the dataset’s utility and highlight substantial performance gaps in challenging conditions, establishing a new platform for advancing depth estimation research. Dataset will be available at https://xiandaguo.net/ROVR-Open-Dataset/
@@ I Introduction
Depth estimation [1, 2] is a cornerstone of 3D scene understanding, providing essential geometric information for a wide range of applications, including autonomous driving [1, 2, 3], robotics [4], and augmented reality (AR) [5]. Accurate depth maps enable downstream tasks such as obstacle detection, motion planning, semantic mapping, and human–machine interaction. Over the past decade, the availability of high-quality annotated datasets has been a major driver of progress in this field, enabling the development of increasingly sophisticated learning-based approaches. Notable examples include large-scale driving datasets such as KITTI [6, 7], nuScenes [8], and DDAD [9], which have served as critical benchmarks for monocular and stereo depth estimation. However, despite their impact, each of these datasets presents limitations in one or more aspects, including scene diversity, annotation density, sensor calibration, and scalability. KITTI [10], while offering high-precision LiDAR depth, is restricted to a relatively narrow set of urban and suburban scenarios, and its limited scale constrains the training of high-capacity models. nuScenes [8] broadens the range of environments and weather conditions but suffers from lower-quality ground truth due to LiDAR calibration inaccuracies. DDAD [9] achieves dense and long-range depth measurements using a sophisticated multi-sensor setup, yet its reliance on costly hardware and complex calibration makes large-scale collection economically challenging.
With the rapid evolution of deep neural networks–particularly transformer-based architectures [1, 11, 12] and multi-modal foundation models–there has been a significant surge in demand for large-scale, diverse, and challenging datasets. As the capacity of models grows, their performance on existing benchmarks is approaching saturation, leaving diminishing room for further improvement without fundamentally new data sources. This trend is especially evident in autonomous driving, where domain-specific datasets can no longer keep pace with the generalization requirements of modern models.
@@
Beyond providing large-scale training data, our dataset introduces new challenges for the community: the driving diversity is significantly higher and differs from those in KITTI or DDAD, requiring models to adapt to varying sampling distributions, demanding stronger generalization ability; meanwhile, the depth GT density is lower than existing benchmarks, which is harder for model to learn. We summarize key comparison against existing datasets in Table I, present scene-wise distributions in Table II, and show representative samples in Figure 2. The capture setup and sensor configuration are illustrated in Figure 1. We further benchmark standard monocular depth estimation baselines on our dataset (Section IV), demonstrating its utility and the remaining gap to close in challenging scenarios. To provide a technical understanding of the proposed ROVR open dataset, we benchmarked several top-ranked KITTI models in Table IV and domain-specific testing in Table V. Notably, KITTI SOTA models are far from saturating depth estimation metrics on ROVR, highlighting its greater difficulty and broader coverage of real-world scenarios. ROVR provides a crucial platform for developing depth estimation models with stronger generalization to complex, dynamic, real-world scenarios.
@@
KITTIDepth [10] Driving 85k 7k 1226 $\times$ 370
nuScenes [8] Driving 34149 5851 1600 $\times$ 900
@@ II Related Work
Other Depth Datasets. A variety of datasets target depth estimation beyond autonomous driving scenarios, covering indoor, outdoor, and mixed environments, as well as synthetic domains. NYUv2 [13] provides 795 training and 654 test frames at $640\times 480$ resolution for indoor RGB-D scenes. ScanNet [14] offers a large-scale collection of 2.5M frames with dense annotations for mixed indoor environments at the same resolution. MegaDepth [15] supplies $\sim$128K internet-photo-derived frames via multi-view stereo, spanning diverse indoor/outdoor scenes. TUM RGB-D [16] focuses on benchmarking depth and camera pose estimation with $640\times 480$ resolution RGB-D sequences. Synthetic datasets such as SceneNet RGB-D [17] (5M training and 300K test frames, $320\times 240$) enable controlled, noise-free depth generation for pre-training and domain adaptation. Driving-oriented datasets like Cityscapes [18], RobotCar [19], and KITTIDepth [10] expand coverage for outdoor urban scenes, with KITTIDepth offering 85K/7K stereo frames with ground-truth depth. More recent large-scale driving datasets include DrivingStereo [20] (174K/7.7K frames, $1762\times 800$), nuScenes [8] (34K/5.8K frames, $1600\times 900$), DDAD [9] (17K/4.1K frames, $1936\times 1216$), and Waymo [21] (600K frames, $1920\times 1280$). Mixed-scene datasets such as DIODE [22] and DA-2K [11] offer diverse environments with dense depth, while our proposed ROVR dataset provides 193K/10K high-resolution ($1920\times 1280$) driving frames with centimeter-level accuracy.
(a) Photo of real vehicle
@@
Figure 1: Illustration of the data collection vehicles: (a)real-world vehicle used for urban field tests; (c) sensor suite on Vehicle 3 including IMU, GNSS/RTK, camera, LiDAR, and computing platform.
Monocular Depth Estimation. The introduction of end-to-end trainable neural networks in MMDE, pioneered by [23], marked a milestone by introducing optimization with the scale-invariant log loss ($\mathrm{SI}_{\log}$). Subsequent advances span convolution-based architectures [24, 25, 26, 27], diffusion-based models [2] and transformer-based models [28, 29, 30, 31, 1], achieving impressive in-domain accuracy—e.g., PackNet [9], DPT [32], and AdaBins [29] report 5–6% relative error on KITTI, approaching LiDAR precision. However, these models often overfit to their training domain and degrade sharply in novel scenes. To improve cross-domain generalization, recent works explore large-scale and diverse training data, as in the Depth-Anything series [12, 33], or adopt generalizable MMDE strategies [34, 35, 36] that incorporate camera awareness [37, 35] or normalize outputs via intrinsics [38, 39, 36]. While geometric pretraining [34] and dataset-specific priors [36] can boost zero-shot accuracy, these methods typically assume noiseless pinhole intrinsics and rely on predefined backprojection, limiting applicability. Benchmarks such as the Monocular Depth Estimation Challenge (MDEC) now track both in-domain and zero-shot performance, with DepthAnything v2 and Marigold [40] as strong recent baselines.
@@ IV Experiments
TABLE V: Depth estimation performance across different conditions. All models are trained on the unified training split and evaluated on different test sub-domains of the same dataset. The upper block reports results under varying illumination (light conditions), while the lower block reports results across scene types.
@@
Domain Method silog abs_rel log10 rms sq_rel log_rms $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
(a) Illumination Conditions
Normal VA-DepthNet [41] 29.8942 0.1833 0.0528 7.0396 4.0945 0.3035 0.8884 0.9366 0.9573
@@
IEBins [44] 30.3786 0.2101 0.0648 7.4493 3.7839 0.3125 0.8501 0.9238 0.9509
Night VA-DepthNet [41] 27.3737 0.1716 0.0493 6.6754 3.4274 0.2770 0.8804 0.9368 0.9617
@@
IEBins [44] 28.3425 0.2081 0.0589 6.8354 3.7938 0.2914 0.8497 0.9263 0.9592
Rainy VA-DepthNet [41] 25.5215 0.1709 0.0538 6.3357 3.0841 0.2597 0.8684 0.9360 0.9609
@@
IEBins [44] 27.6484 0.2891 0.1025 7.3199 3.2361 0.3233 0.7015 0.9163 0.9532
(b) Scene Types
Highway VA-DepthNet [41] 26.2820 0.1476 0.0431 6.2185 3.1248 0.2661 0.9106 0.9484 0.9663
IEBins [44] 27.0938 0.1809 0.0565 6.6169 3.0320 0.2807 0.8690 0.9373 0.9613
Rural VA-DepthNet [41] 29.1695 0.1803 0.0489 6.9052 4.1370 0.2948 0.8980 0.9420 0.9596
IEBins [44] 29.9301 0.2372 0.0731 7.4677 3.8916 0.3211 0.8352 0.9312 0.9550
Urban VA-DepthNet [41] 38.0108 0.2763 0.0802 9.0505 6.2298 0.3883 0.8131 0.9013 0.9341
IEBins [44] 38.1213 0.3012 0.0916 9.4542 5.7344 0.3911 0.7645 0.8823 0.9246
@@ IV-A Implementation Details
We evaluate three representative monocular depth baselines on our dataset—VA-DepthNet [41], DCDepth [43], and IEBins [44]. To ensure comparability, we reproduce the variants that adopt the Swin Transformer–Large (Swin-L) [45] backbone and initialize training solely from its pretrained weights. All images are center-cropped to $960\times 544$ for both training and evaluation, and each model is trained for $5$ epochs. Unless otherwise noted, we follow the official training configurations of each method (e.g., learning rate, weight decay). Ground-truth depth is measured in meters; during training and evaluation, depths are clipped to $[1,80]$ meters and invalid pixels are masked. All experiments are conducted on $8\times$ NVIDIA L20 GPUs.

### v2

@@ Abstract
Depth estimation is a fundamental task for 3D scene understanding in autonomous driving, robotics, and augmented reality. Existing depth datasets, such as KITTI, nuScenes, and DDAD, have advanced the field but suffer from limitations in diversity and scalability. As benchmark performance on these datasets approaches saturation, there is an increasing need for a new generation of large-scale, diverse, and cost-efficient datasets to support the era of foundation models and multi-modal learning. We present ROVR, a large-scale, diverse, and cost-efficient depth dataset designed to capture the complexity of real-world driving. ROVR comprises 200K high-resolution frames across highway, rural, and urban scenarios, spanning day/night and adverse weather conditions. A lightweight acquisition pipeline ensures scalable collection, while sparse but statistically sufficient ground truth supports robust training. Benchmarking with state-of-the-art monocular depth models reveals severe cross-dataset generalization failures: models achieving near-ceiling accuracy on KITTI degrade drastically on ROVR, and even when trained on ROVR, current methods fall short of saturation. These results highlight the unique challenges posed by ROVR—scene diversity, dynamic environments, and sparse ground truth—establishing it as a demanding new platform for advancing depth estimation and building models with stronger real-world robustness. Extensive ablation studies provide a more intuitive understanding of our dataset across different scenarios, lighting conditions, and generalized ability. Dataset will be available at https://xiandaguo.net/ROVR-Open-Dataset/
@@ I Introduction
Depth estimation is a cornerstone of 3D scene understanding, providing essential geometric information for a wide range of applications, including autonomous driving [1, 2], robotics [3, 4], and augmented reality (AR) [5]. Accurate depth maps enable downstream tasks such as obstacle detection, motion planning, semantic mapping, and human–machine interaction. Over the past decade, the availability of high-quality annotated datasets has been a major driver of progress in this field, enabling the development of increasingly sophisticated learning-based approaches. Notable examples include large-scale driving datasets such as KITTI [6, 7], nuScenes [8], and DDAD [9], which have served as critical benchmarks for monocular and stereo depth estimation. However, despite their impact, each of these datasets presents limitations in one or more aspects, including scene diversity, annotation density, sensor calibration, and scalability. KITTI [10], while offering high-precision LiDAR depth, is restricted to a relatively narrow set of urban and suburban scenarios, and its limited scale constrains the training of high-capacity models. nuScenes [8] broadens the range of environments and weather conditions but suffers from lower-quality ground truth due to LiDAR calibration inaccuracies. DDAD [9] achieves dense and long-range depth measurements using a sophisticated multi-sensor setup, yet its reliance on costly hardware and complex calibration makes large-scale collection economically challenging.
With the rapid evolution of deep neural networks–particularly transformer-based architectures [11, 12] and multi-modal foundation models–there has been a significant surge in demand for large-scale, diverse, and challenging datasets. As the capacity of models grows, their performance on existing benchmarks is approaching saturation, leaving diminishing room for further improvement without fundamentally new data sources. This trend is especially evident in autonomous driving, where domain-specific datasets can no longer keep pace with the generalization requirements of modern models.
@@
Beyond providing large-scale training data, our dataset introduces new challenges for the community: the driving diversity is significantly higher and differs from those in KITTI or DDAD, requiring models to adapt to varying sampling distributions, demanding stronger generalization ability; meanwhile, the depth ground truth density is lower than existing benchmarks, which is harder for model to learn. We summarize key comparison against existing datasets in Table I, present scene-wise distributions in Table II, and show representative samples in Figure 2. The capture setup and sensor configuration are illustrated in Figure 1. We further benchmark standard monocular depth estimation baselines on our dataset (Section IV), demonstrating its utility and the remaining gap to close in challenging scenarios. To provide a technical understanding of the proposed ROVR open dataset, we benchmarked several top-ranked KITTI models in Table IV and domain-specific testing in Table VVI. Notably, SOTA models on KITTI are far from saturating depth estimation metrics on ROVR, highlighting its greater difficulty and broader coverage of real-world scenarios. ROVR provides a crucial platform for developing depth estimation models with stronger generalization to complex, dynamic, real-world scenarios.
@@
KITTIDepth [10] Driving 86,000 7,000 1226 $\times$ 370
nuScenes [8] Driving 34,149 5,851 1600 $\times$ 900
@@ II Related Work
Other Depth Datasets. A variety of datasets target depth estimation beyond autonomous driving scenarios, covering indoor, outdoor, and mixed environments, as well as synthetic domains. NYUv2 [13] provides 795 training and 654 test frames at $640\times 480$ resolution for indoor RGB-D scenes. ScanNet [14] offers a large-scale collection of 2.5M frames with dense annotations for mixed indoor environments at the same resolution. MegaDepth [15] supplies $\sim$128K internet-photo-derived frames via multi-view stereo, spanning diverse indoor/outdoor scenes. TUM RGB-D [16] focuses on benchmarking depth and camera pose estimation with $640\times 480$ resolution RGB-D sequences. Synthetic datasets such as SceneNet RGB-D [17] (5M training and 300K test frames, $320\times 240$) enable controlled, noise-free depth generation for pre-training and domain adaptation. Driving-oriented datasets like Cityscapes [18], RobotCar [19], and KITTIDepth [10] expand coverage for outdoor urban scenes, with KITTIDepth offering 85K/7K stereo frames with ground-truth depth. More recent large-scale driving datasets include DrivingStereo [20] (174K/7.7K frames, $1762\times 800$), nuScenes [8] (34K/5.8K frames, $1600\times 900$), DDAD [9] (17K/4.1K frames, $1936\times 1216$), and Waymo [21] (600K frames, $1920\times 1280$). Mixed-scene datasets such as DIODE [22] and DA-2K [23] offer diverse environments with dense depth, while our proposed ROVR dataset provides 193K/10K high-resolution ($1920\times 1280$) driving frames with centimeter-level accuracy. Finally, the Booster dataset [24, 25] covers indoor scenes featuring transparent and reflective objects.
(a) Real vehicle
@@
Figure 1: Illustration of the data collection vehicles: (a)real-world vehicle used for urban field tests; (c) sensor suite on Vehicle 3, including IMU, GNSS/RTK, camera, LiDAR, and computing platform.
Monocular Depth Estimation. The introduction of end-to-end trainable neural networks in MMDE, pioneered by [26], marked a milestone by introducing optimization with the scale-invariant log loss ($\mathrm{SI}_{\log}$). Subsequent advances span convolution-based architectures [27, 28, 29, 30], diffusion-based models [1] and transformer-based models [31, 32, 33, 34, 11], achieving impressive in-domain accuracy—e.g., PackNet [9], DPT [35], and AdaBins [32] report 5–6% relative error on KITTI, approaching LiDAR precision. However, these models often overfit to their training domain and degrade sharply in novel scenes. To improve cross-domain generalization, recent works explore large-scale and diverse training data, as in the Depth-Anything series [36, 23], or adopt generalizable MMDE strategies [37, 38, 39] that incorporate camera awareness [40, 38] or normalize outputs via intrinsics [41, 42, 39]. While geometric pretraining [37] and dataset-specific priors [39] can boost zero-shot accuracy, these methods typically assume noiseless pinhole intrinsics and rely on predefined backprojection, limiting applicability. Benchmarks such as the Monocular Depth Estimation Challenge (MDEC) [43, 44, 45, 46] now track both in-domain and zero-shot performance, with DepthAnything v2 and Marigold [47] as strong recent baselines.
@@ IV Experiments
TABLE V: Depth estimation performance under varying illumination (light conditions). Zero-shot methods are divided into affine-invariant (top) and metric (bottom).
@@
Domain Model Type Method silog abs_rel log10 rms sq_rel log_rms $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
Normal Marigold v1.1 [47] 59.6965 0.9370 0.2595 14.2185 15.1616 0.7109 0.1993 0.3969 0.6006
DepthAnythingV2 [23] 71.4943 0.9399 0.2721 14.4564 15.3124 0.8032 0.1974 0.3933 0.5955
DepthAnythingV2 [23] 59.5318 0.9390 0.2595 14.2243 15.2125 0.7105 0.1992 0.3968 0.6005
Depth Pro [52] 59.5396 0.9367 0.2592 14.2148 15.1633 0.7098 0.1996 0.3975 0.6013
UniK3D [53] 59.5338 0.9390 0.2595 14.2244 15.2125 0.7105 0.1992 0.3968 0.6005
Zero-Shot UniDepth v2 [54] 59.5325 0.9390 0.2595 14.2243 15.2125 0.7105 0.1992 0.3968 0.6005
VA-DepthNet [48] 29.8942 0.1833 0.0528 7.0396 4.0945 0.3035 0.8884 0.9366 0.9573
@@
Single-Domain IEBins [51] 30.3786 0.2101 0.0648 7.4493 3.7839 0.3125 0.8501 0.9238 0.9509
Night Marigold v1.1 [47] 54.2200 0.9902 0.2771 13.6023 14.9837 0.7303 0.1520 0.3255 0.5256
DepthAnythingV2 [23] 58.0942 0.9896 0.2804 13.6943 15.0012 0.7542 0.1517 0.3248 0.5245
DepthAnythingV2 [23] 54.2149 0.9902 0.2771 13.6024 14.9842 0.7302 0.1520 0.3255 0.5256
Depth Pro [52] 54.2161 0.9902 0.2771 13.6025 14.9838 0.7302 0.1520 0.3255 0.5256
UniK3D [53] 54.2150 0.9902 0.2771 13.6024 14.9842 0.7302 0.1520 0.3255 0.5256
Zero-Shot UniDepth v2 [54] 54.2149 0.9902 0.2771 13.6024 14.9842 0.7302 0.1520 0.3255 0.5256
VA-DepthNet [48] 27.3737 0.1716 0.0493 6.6754 3.4274 0.2770 0.8804 0.9368 0.9617
@@
Single-Domain IEBins [51] 28.3425 0.2081 0.0589 6.8354 3.7938 0.2914 0.8497 0.9263 0.9592
Rainy Marigold v1.1 [47] 57.9063 1.2563 0.3158 14.6217 20.9788 0.8416 0.1607 0.3243 0.4829
DepthAnythingV2 [23] 69.7717 1.2578 0.3286 14.7804 21.0635 0.9267 0.1586 0.3202 0.4772
DepthAnythingV2 [23] 57.8361 1.2563 0.3157 14.6209 20.9781 0.8412 0.1607 0.3243 0.4830
Depth Pro [52] 57.8361 1.2563 0.3157 14.6209 20.9781 0.8412 0.1607 0.3243 0.4830
UniK3D [53] 57.8363 1.2563 0.3157 14.6209 20.9781 0.8412 0.1607 0.3243 0.4830
Zero-Shot UniDepth v2 [54] 57.8361 1.2563 0.3157 14.6209 20.9781 0.8412 0.1607 0.3243 0.4830
VA-DepthNet [48] 25.5215 0.1709 0.0538 6.3357 3.0841 0.2597 0.8684 0.9360 0.9609
@@
Single-Domain IEBins [51] 27.6484 0.2891 0.1025 7.3199 3.2361 0.3233 0.7015 0.9163 0.9532
@@ IV-A Implementation Details
We evaluate two families of depth estimation baselines.
Zero-Shot Models. Networks belonging to this category have been trained on web-scale data, either to predict affine-invariant or metric depth, to excel at zero-shot generalization on unseen domains. We select two affine-invariant and four metric baselines: specifically, Marigold v1.1 [47] and DepthAnythingV2 [23] for the former group, KITTI fine-tuned DepthAnythingV2 [23], Depth Pro [52], UniK3D [53], and UniDepth v2 [54] for the latter. We directly use the weights provided by the authors, selecting those yielding the most accurate results for each model.
Single-domain Models. Networks belonging to this category are usually trained to estimate metric depth on a single large-scale dataset, aiming at excelling in a specific domain. We select three representative baselines: VA-DepthNet [48], DCDepth [50], and IEBins [51]. To ensure comparability, we reproduce the variants that adopt the Swin Transformer–Large (Swin-L) [55] backbone and initialize training solely from its pretrained weights. All images are center-cropped to $960\times 544$ for both training and evaluation, and each model is trained for $5$ epochs. Unless otherwise noted, we follow the official training configurations of each method (e.g., learning rate, weight decay). Ground-truth depth is measured in meters; during training and evaluation, depths are clipped to $[1,80]$ meters and invalid pixels are masked. All experiments are conducted on $8\times$ NVIDIA L20 GPUs.
[…]
@@ IV-B Evaluation Metrics
TABLE VI: Depth estimation performance across scene types. Zero-shot methods are divided into affine-invariant (top) and metric (bottom).
[…]
Domain Model Type Method silog abs_rel log10 rms sq_rel log_rms $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
Highway Marigold v1.1 [47] 59.2410 0.9960 0.2719 14.2594 16.0318 0.7353 0.1784 0.3640 0.5659
DepthAnythingV2 [23] 69.1255 0.9982 0.2821 14.4595 16.1547 0.8102 0.1771 0.3614 0.5620
DepthAnythingV2 [23] 59.0573 0.9983 0.2718 14.2660 16.0903 0.7348 0.1784 0.3640 0.5659
Depth Pro [52] 59.0670 0.9955 0.2715 14.2549 16.0308 0.7341 0.1787 0.3648 0.5668
UniK3D [53] 59.0593 0.9983 0.2718 14.2662 16.0903 0.7348 0.1784 0.3640 0.5659
Zero-Shot UniDepth v2 [54] 59.0581 0.9983 0.2718 14.2661 16.0903 0.7348 0.1784 0.3640 0.5659
VA-DepthNet [48] 26.2820 0.1476 0.0431 6.2185 3.1248 0.2661 0.9106 0.9484 0.9663
Single-Domain IEBins [51] 27.0938 0.1809 0.0565 6.6169 3.0320 0.2807 0.8690 0.9373 0.9613
Rural Marigold v1.1 [47] 56.1643 0.9741 0.2653 13.8753 15.6193 0.7193 0.2016 0.4003 0.5931
DepthAnythingV2 [23] 69.0038 0.9761 0.2786 14.1191 15.7402 0.8198 0.1992 0.3959 0.5872
DepthAnythingV2 [23] 56.1199 0.9741 0.2653 13.8749 15.6189 0.7190 0.2016 0.4003 0.5931
Depth Pro [52] 56.1199 0.9741 0.2653 13.8749 15.6189 0.7190 0.2016 0.4003 0.5931
UniK3D [53] 56.1207 0.9741 0.2653 13.8749 15.6189 0.7190 0.2016 0.4003 0.5931
Zero-Shot UniDepth v2 [54] 56.1199 0.9741 0.2653 13.8749 15.6189 0.7190 0.2016 0.4003 0.5931
VA-DepthNet [48] 29.1695 0.1803 0.0489 6.9052 4.1370 0.2948 0.8980 0.9420 0.9596
Single-Domain IEBins [51] 29.9301 0.2372 0.0731 7.4677 3.8916 0.3211 0.8352 0.9312 0.9550
Urban DepthAnythingV2 [23] 71.9928 0.8663 0.2598 14.2754 14.0632 0.7837 0.2214 0.4299 0.6280
DepthAnythingV2 [23] 59.0757 0.8636 0.2454 14.0271 13.9164 0.6812 0.2240 0.4345 0.6341
Depth Pro [52] 59.0758 0.8635 0.2454 14.0259 13.9145 0.6812 0.2240 0.4346 0.6341
Marigold v1.1 [47] 59.1288 0.8632 0.2454 14.0258 13.9058 0.6815 0.2240 0.4344 0.6340
UniK3D [53] 59.0768 0.8636 0.2454 14.0272 13.9164 0.6812 0.2240 0.4344 0.6341
Zero-Shot UniDepth v2 [54] 59.0760 0.8636 0.2454 14.0271 13.9164 0.6812 0.2240 0.4344 0.6341
VA-DepthNet [48] 38.0108 0.2763 0.0802 9.0505 6.2298 0.3883 0.8131 0.9013 0.9341
Single-Domain IEBins [51] 38.1213 0.3012 0.0916 9.4542 5.7344 0.3911 0.7645 0.8823 0.9246